How much energy does a photon of blue light with a wavelength of 465 nm have?

How much energy does a photon of red light with a wavelength of 663 nm have?
Photons of which color of light (red or blue) have greater energy?

red or blue    
Answer #1


     = 465nm   = 465*10^-9 m

C   = 3*10^8m/s

E   = hc/

     = 6.625*10^-34*3*10^8/465*10^-9

      = 4.274*10^-19 J of blue light

     = 663nm   = 663*10^-9 m

C   = 3*10^8m/s

E   = hc/

     = 6.625*10^-34*3*10^8/663*10^-9

       = 3*10^-19 J of red light

blue light has greater energy than red light

blue >>>>answer